OpenStudy (jkbo):

What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form.

OpenStudy (anonymous):

y = 4x

OpenStudy (anonymous):

Well, we have\[y=mx + b\]\[m = (y _{1}-y _{2})/(x _{1}-x _{2})\]\[m=(-4-4)/(-1-1)=4\]So,\[y=4x+b\]Taking point (-1,-4) into account we have,\[-4=4\times(-1)+b \rightarrow -4=-4+b \rightarrow b=0\] Therefor,\[y=4x\]